### v1

@@ I Introduction
This study aims to contribute to preserving historic buildings by creating digital twins. Unlike traditional 3D models commonly used in building information modeling (BIM), parametric models were employed to capture essential attributes and behaviors of buildings. The main contributions of this work are:
  • Presented a cloud-based solution to create parametric digital twins for conservation of historic buildings by combining graphical and relational data models. The graphical model stores contextual data, such as spatial information of rooms and sensor locations. In contrast, the relational model captures time series data like indoor climate and outdoor weather conditions. A full implementation of this approach was also given. The implementation employed and extended the Brick ontology [21] to ensure data consistency and interoperability, enhancing the integration of diverse data sources and applications. To avoid reinvention, several public cloud services provided by Microsoft Azure [23] were utilized to build the cloud part of the solution.
  • Conducted a comprehensive case study in a historic building in Östergötland, Sweden. Thirteen cloud-connected sensor boxes, equipped with 84 sensors, were deployed across all floors to create a parametric digital twin of the building. Data analytics were then performed to evaluate indoor climate in the building. The results revealed the necessity of implementing measures to address high humidity problems in the basement and on the ground floor, e.g., installing vapor barriers. Meanwhile, energy-efficient heating and ventilation strategies were identified for the upper floors. Notably, the presented solution and insights derived from this study are transferable to other historic buildings facing similar problems.
@@ III-B Implementation
Our implementation is a comprehensive solution for creating parametric digital twins of historic buildings. It integrates multiple environmental sensors for collecting data about indoor climate, utilizes an edge platform for communication, combines graphical and relational models for organizing data, and leverages Microsoft Azure cloud services for data storage and deployment of developed applications. We have also included provisions for future expansion of the implementation.
@@ III-B1 Edge Platform and Sensors
  • PPD42NS (Shinyei Corporation, New York, USA) detects particulate matter with a diameter larger than 1 $\mu$m.
@@
As shown in Fig. 2, the edge platform and six sensors are packaged in a plastic box (size: 201$\times$163$\times$98 mm). More sensors can be added based on specific need. For example, in the case study described in Section IV, two additional ultrasonic sensors (UR18.DA0.2-UAMJ.9SF, Baumer, Frauenfeld, CHE) were added to the sensor box to monitor groundwater level changes in the basement.
@@ IV-A Description of Löfstad Castle
Löfstad Castle is located about 9 km southwest of the city of Norrköping in Östergötland, Sweden. The main building of the castle (see Fig. 3) has a basement, three floors above the basement (ground, first, and second floors), and an attic. The entrance of the main building faces east and towards a courtyard flanked by two wings on the north and south sides. The main building was constructed with heavy masonry, with outer walls approximately 1.1 m thick. Currently, the main building is naturally ventilated, and only the first floor has electric radiant heating during the winter season.
@@
Löfstad Castle holds significant cultural and historic values. It has been a designated historic monument since 1983 [59]. The castle was first built in the 17th century. In 1750, the castle was burned down due to a great fire. Both the main building and the north wing were severely damaged by fire. Apart from a few objects from the main building, the south wing was saved. The castle was then rebuilt and got its current form with a baroque characteristic. Miss Emilie Piper was the last owner of Löfstad Castle [60]. Before Emilie died in 1926, she donated the inventory to Östergötland’s Museum with clear instructions to preserve the interior of the stateroom on the first floor exactly as it was. The estate property is managed by Riddarhuset through a foundation. The Löfstad Castle Museum has been open to the public since 1942. Over the years, only minimal changes have occurred to the interior of the stateroom and the collection it houses. Other parts of the interior of the main building have slightly changed, but its historic character has been mostly retained [61]. These efforts make the castle a remarkably well-preserved example of manor buildings from the 17th and 18th centuries [62].
However, the historic interior of the main building and housed collections are deteriorating due to inappropriate indoor climate, primarily high RH. The problem originates in the basement, where the earthen floor allows groundwater to evaporate from soil into the air continuously. This moisture is speculated to rise through walls, increasing humidity levels on upper floors. High RH poses risks of mold and other biological damage. The high humidity creates ideal conditions for woodworm, which causes considerable harm to wooden objects in the building.
@@ IV-B Deployment of Local Devices
The case study focuses on creating a parametric digital twin to help users better understand indoor climate in the main building. As shown in Fig. 4, a total of 84 sensors are distributed throughout the building, from the basement to the attic, to continuously monitor indoor environmental conditions on all floors. The number of sensor boxes deployed in the basement (BF), on the ground floor (GF), first floor (1F), second floor (2F), and attic is three, three, three, three, and one, respectively, with each sensor box containing sensors as described in Section III-B1.
@@
Placements of sensor boxes and groundwater level (GWL) sensors are indicated on floor plans of the main building (see Fig. 5). Details on the location, height above the floor, and installation date of sensor boxes are provided in Table I. All six GWL sensors were also installed on June 28, 2023. The positioning of sensor boxes was carefully chosen to meet both building conservation requirements and site conditions. Examples of a deployed sensor box and a GWL sensor are illustrated in Fig. 6. In Room 103 (the stateroom) on the first floor, the sensor box is placed on top of a cabinet (see Fig. 6a). Fig. 6b shows a GWL sensor, which uses ultrasound to measure the distance from the water surface to the probe.
@@
All sensors were calibrated prior to deployment. Sensors for measuring temperature, RH, and GWL were calibrated by ourselves. The remaining sensors were calibrated by manufacturers. Each environmental parameter is collected every 30 seconds to capture rapid fluctuations, with the option for downsampling during subsequent data analytics and visualization.
@@ IV-C Collection of External Data
Collected external data are outdoor weather conditions that affect indoor climate, including dry-bulb temperature, RH, dew point temperature, precipitation, and air pressure. Hourly measurements of these parameters were obtained through open APIs provided by the Swedish Meteorological and Hydrological Institute [63]. The data were measured from a weather station (ID: 86340) located about 7 km from Löfstad Castle. These measurements serve as an approximation of ambient weather conditions surrounding the building.
@@ IV-D Developed Applications
Four data applications were developed, focusing on data visualization, analytics, and sharing. Fig. 7a presents the application to visualize time plots of environmental parameters collected from a selected sensor box within a specified date range. This tool helps uers inspect individual monitoring points, identify trends and fluctuations, as well as explore correlations between different environmental parameters. For example, based on the one-week data shown in Fig. 7a, it is evident that temperature exhibited strong daily seasonality, while CO2 concentration showed six distinct peaks, indicating different levels of occupancy in Room 103 on those days.
@@
Open data access allows project participants and external researchers to obtain data for further analysis from various perspectives. The data sharing application (available at https://historicbuildings.azurewebsites.net/download) enables users to choose various sampling intervals (raw or downsampled), parameters, and time periods. Downloaded data are cleaned of anomalies, such as out-of-range values, and provided in comma-separated values (CSV) format, making it accessible for analysis with various tools.
@@ IV-E Analysis of Indoor Climate
For question 2, indoor and outdoor humidity mixing ratios (MRs) over a year were compared to identify potential moisture sources. The MR, expressed in g/kg, is the ratio of the mass of water vapor to the mass of dry air [26]. A comparison between indoor and outdoor MRs suggests a water source or sink in a building [27]. If the average indoor MR consistently exceeds the outdoor MR, it might indicate constant evaporation from the ground or walls of the building. MR is calculated by (1).
[…]
where $t$ is temperature (C), $RH$ is relative humidity (%), $p$ is atmospheric pressure (hPa). Since indoor atmospheric pressure was not measured, $p$ was substituted with a standard pressure of 1013 hPa for approximate calculations of indoor MRs, as recommended by the standard EN 16242:2012 [26]. Outdoor MRs were calculated using collected atmospheric pressure.
@@
$LIM_{I}=\ cosh(0.128324\times(30-t))+75,$ (2)
where $t$ is temperature (C). The calculated value is expressed as a percentage (%). A RH value exceeding the LIMI at a given temperature is considered risky. Short-term fluctuations of RH throughout a year were examined to determine the target range of RH and identify risks associated with strain-stress cycles, following the standard EN 15757:2010 [24].
@@ V-A Temperature and Relative Humidity Throughout a Year
In rooms not equipped with heating systems, the trend of indoor temperature closely follows outdoor temperature. In the main building of Löfstad Castle, rooms in the basement and rooms on the ground and second floors lack heating systems. As shown in Fig. 8a and Fig. 8b, temperature in these rooms exhibited a similar trend to outdoor temperature throughout the year. Rooms on the first floor are equipped with heating systems. When these systems were turned off, the trend of indoor temperature was also driven by outdoor temperature. This pattern was evident from May to September 2023 (see Fig. 8b) for rooms on the first floor. However, when heating systems were working during the heating season, these rooms maintained a stable temperature around 16 C.
@@
Indoor temperature of different rooms on the same floor tend to be relatively close, while noticeable differences occur between rooms on different floors. During summer months, indoor temperature generally increases as a floor goes up. As depicted in Fig. 8b, around early July 2023, temperature in rooms on the first and second floors was approximately 25 C. Temperature of Room 222 on the second floor even reached up to 30 C. In contrast, during the same period, temperature of rooms on the ground floor and in the basement were approximately 20 C and 12 C, respectively (see Fig. 8a).
Abnormal variations in indoor temperature were observed during specific periods. As shown in Fig. 8a, Room 17 on the ground floor experienced an unusually high temperature in the first half of December 2023. Moreover, heated rooms like Rooms 113 and 117 exhibited temperature spikes in November and December 2023 (see Fig. 8b). These anomalies may be attributed to maintenance activities conducted in these rooms.
@@
Rooms on the same floor typically have similar levels of RH. As depicted in Fig. 9a, RH in both basement rooms remained close to 100% throughout the year. This extremely high humidity is primarily due to the earthen floor, which allows groundwater to continuously evaporate from the wet soil into the indoor air. Likewise, Room 3 and Room 17 on the ground floor (see Fig. 9a), as well as the three rooms on both the first and second floors (see Fig. 9b), each have approximate levels of RH. An exception was observed in Room 5 on the ground floor (see Fig. 9a), where the sensor box detected higher RH compared to other rooms on the same floor. The RH frequently exceeded 80% and approached 100% from July to October 2023. The high RH readings are likely caused by the evaporation of water from the ground and moisture from the basement rise through walls.
There are notable differences in RH across floors. As illustrated in Fig. 9a and Fig. 9b, in rooms without heating systems (BF, GF, and 2F), RH decreased as the floor goes up. Conversely, in rooms with heating systems (1F), RH was lower compared to rooms on the second floor when the heating system was working.

### v2

@@ I Introduction
This study aims to contribute to preserving historic buildings by creating digital twins that extend beyond the traditional 3D models commonly used in building information modeling (BIM). Conventional BIM approaches mainly handle static attributes. This work employed parametric models to capture both essential static attributes and dynamic behaviors of buildings, which enhances more comprehensive environmental monitoring and analysis. The main contributions of this work are:
  • Presented a cloud-based solution to create parametric digital twins for conservation of historic buildings by combining graphical and relational data models. The graphical model stores contextual data, such as spatial information of rooms and sensor locations. In contrast, the relational model captures time series data like indoor climate and outdoor weather conditions. This integration facilitates real-time monitoring and provides data-driven insights into building operations. A full implementation of this approach was also given. The implementation employed and extended the Brick ontology [21] to ensure data consistency and interoperability, enhancing the integration of diverse data sources and applications. To avoid reinvention, several public cloud services provided by Microsoft Azure [23] were utilized to build the cloud part of the solution.
  • Conducted a comprehensive case study in a historic building in Östergötland, Sweden. Thirteen cloud-connected sensor boxes, equipped with 84 sensors, were deployed across all floors. This extensive sensor distribution enabled the creation of a parametric digital twin that not only captured environmental conditions but also supported advanced data analytics to assess the indoor climate. The analysis revealed the necessity of implementing measures to address high humidity problems in the basement and on the ground floor, e.g., installing vapor barriers. Meanwhile, energy-efficient heating and ventilation strategies were identified for the upper floors. Notably, the presented solution and insights derived from this study are transferable to other historic buildings facing similar problems, providing a flexible framework that can be adapted to various environments and conservation challenges.
@@ III-B Implementation
The proposed implementation offers a systematic approach for creating parametric digital twins of historic buildings. It integrates multiple environmental sensors for collecting data about indoor climate, utilizes an edge platform for communication, combines graphical and relational models for organizing data, and leverages Microsoft Azure cloud services for data storage and deployment of developed applications. The implementation also considers scalability, which allows for future expansion and integration of additional features.
@@ III-B1 Edge Platform and Sensors
  • PPD42NS (Shinyei Corporation, New York, USA) detects particulate matter with a diameter larger than 1 µm.
@@
As shown in Fig. 2, the edge platform and six sensors are packaged in a plastic box (size: 201 mm×$\times$163 mm×$\times$98 mm). More sensors can be added based on specific need. For example, in the case study described in Section IV, two additional ultrasonic sensors (UR18.DA0.2-UAMJ.9SF, Baumer, Frauenfeld, CHE) were added to the sensor box to monitor groundwater level changes in the basement.
@@
The performance of these sensors in accuracy and reliability is enough for this work, but the limitations should be considered. The AHT20 sensor has reduced accuracy ($\pm$3%) for measuring RH when the environment has an RH outside the typical range (20–80%). The MH-Z16 sensor has declined accuracy at higher CO2 concentrations ($\pm$(50 ppm $+$ 5% of the reading)). The PPD42NS sensor detects particulate matter (PM) larger than 1 µm but does not distinguish them, making it unsuitable for PM2.5 or PM10 detection. Lastly, the LM358 sensor provides relative noise levels without precise decibel measurements.
@@ IV-A Description of Löfstad Castle
Löfstad Castle is located about 9 km southwest of the city of Norrköping in Östergötland, Sweden. The main building of the castle (see Fig. 3) has a basement, three floors above the basement (ground, first, and second floors), and an attic. The entrance of the main building faces east, towards a courtyard flanked by two wings on the north and south sides. The main building was constructed with heavy masonry, with outer walls approximately 1.1 m thick. Currently, the main building is naturally ventilated, and only the first floor has electric radiant heating during the winter season.
@@
Löfstad Castle holds significant cultural and historic values. It has been a designated historic monument since 1983 [59]. The castle was first built in the 17th century. In 1750, a devastating fire broke out. Both the main building and the north wing were severely damaged while the south wing was largely unaffected. The castle was then rebuilt and got its current form with a Baroque characteristic and some Rococo elements [60]. Miss Emilie Piper (1857–1926) was the last owner of Löfstad Castle [61]. In her will (1923), she donated the majority of the collections in the castle to Östergötland’s Museum, and the estate property to Riddarhuset. She also stated that the interior of the stateroom on the first floor should be preserved exactly as it was. Today, the estate property is managed by Riddarhuset through a foundation. The Löfstad Castle Museum has been open to the public since 1942. Over the years, only minimal changes have occurred to the interior of the stateroom and the collection it houses. Other parts of the interior of the main building have slightly changed, but its historic character has been mostly retained [62]. These efforts make the castle a well-preserved example of manor buildings from the 17th and 18th centuries [61].
However, the historic interior of the main building and housed collections are deteriorating due to inappropriate indoor climate, primarily high RH. The problem originates in the basement, where the floor consists of exposed soil, allowing groundwater to evaporate from the soil into the air continuously. This moisture is speculated to rise through walls, increasing humidity levels on upper floors. High RH poses risks of mold and other biological damage. The high humidity creates ideal conditions for woodworm, which causes considerable harm to wooden objects in the building (see Fig. 4).
Figure 4: Examples of wooden objects damaged by woodworms: (a) beam of a loom and (b) a wooden ladle.
@@ IV-B Deployment of Local Devices
As shown in Fig. 5, a total of 84 sensors are distributed throughout the building, from the basement to the attic, to continuously monitor indoor environmental conditions on all floors. The number of sensor boxes deployed in the basement (BF), on the ground floor (GF), first floor (1F), second floor (2F), and attic is three, three, three, three, and one, respectively, with each sensor box containing sensors as described in Section III-B1.
@@
Placements of sensor boxes and groundwater level (GWL) sensors are indicated on floor plans of the main building (see Fig. 6). Details on the location, height above the floor, and installation date of sensor boxes are provided in Table I. All six GWL sensors were also installed on June 28, 2023. The positioning of sensor boxes was carefully chosen to meet both building conservation requirements and site conditions. The considered criteria include:
  • Strategic coverage: Sensor boxes were primarily distributed in rooms containing valuable collections, with a secondary emphasis on ensuring even coverage across rooms on different floors. This helps to understand how environmental conditions vary throughout the building.
  • Minimizing interference: Sensor boxes were deliberately positioned away from heat sources, such as radiators, to avoid measurement bias.
  • Monitoring vulnerable areas: Special attention was given to locations susceptible to environmental fluctuations, including areas near windows, exterior walls, and doors, to effectively capture changes in temperature and humidity.
  • Height considerations: In the absence of other considerations, sensor boxes were placed at approximately eye level (1.5–2 m) to avoid skewed measurements.
  • Non-intrusive installation: To preserve structural and historical integrity of the building, sensor boxes were installed without causing damage to walls and floors.
@@
The data continuously collected by these sensors are essential for creating a parametric digital twin and evaluating indoor climate of the building. Specifically, the sensors capture dynamic environmental parameters, which are then integrated into the digital twin to create a detailed and constantly updated model of the building. This enables real-time monitoring and provides a data-driven approach to support proactive decision-making for historic building conservation.
Examples of a deployed sensor box and a GWL sensor are illustrated in Fig. 7. In Room 103 (the stateroom) on the first floor, the sensor box is placed on top of a cabinet (see Fig. 7a). This sensor box provides insights into both indoor climate and occupancy pattern of the room. Fig. 7b shows a GWL sensor, which uses ultrasound to measure the distance from the water surface to the probe. These measurements of GWL are integrated into the parametric digital twin, where they are analyzed alongside outdoor precipitation to identify correlations between rainfall and groundwater levels. Understanding this relationship helps to evaluate potential risks such as rising dampness. This insight facilitates implementing proactive conservation strategies within the digital twin framework, such as optimizing humidity control, to preserve the building.
@@
All sensors were calibrated before deployment. Sensors for measuring temperature, RH, and GWL were calibrated by ourselves. The remaining sensors were calibrated by manufacturers. Each environmental parameter is collected every 30 seconds (two samples per minute) to capture rapid fluctuations, with the possibility for downsampling during subsequent data analytics and visualization.
@@ IV-C Collection of External Data
Collected external data are outdoor weather conditions that affect indoor climate, including dry-bulb temperature, RH, dew point temperature, precipitation, and air pressure. Hourly measurements of these parameters were obtained through open APIs provided by the Swedish Meteorological and Hydrological Institute [63]. The data were measured from a weather station (ID: 86340) located about 7 km from Löfstad Castle. These measurements serve as an approximation of ambient weather conditions surrounding the building. The data obtained from the weather station were considered sufficient for this work since the deviations were negligible. While deploying environmental sensors outside the building would provide more localized and accurate ambient weather data, it should fulfill constraints for historic building conservation to minimize any external alterations that could affect its aesthetics or structure.
@@ IV-D Developed Applications
Four data applications were developed, focusing on data visualization, analytics, and sharing. Fig. 8a presents the application to visualize time plots of environmental parameters collected from a selected sensor box within a specified date range. This tool helps users inspect individual monitoring points, identify trends and fluctuations, as well as explore correlations between different environmental parameters. For example, based on the one-week data shown in Fig. 8a, it is evident that temperature exhibited strong daily seasonality, while CO2 concentration showed six distinct peaks, indicating different levels of occupancy in Room 103 on those days. Moreover, during each day of the five days from July 4 to 8, a noticeable drop followed by an increase in temperature was observed around midday (see Fig. 8a). This fluctuation is speculated to be associated with human activity, especially a temporary leave followed by a return. Pearson correlation coefficient between temperature and CO2 concentration over the five days is 0.64, indicating a strong positive correlation [64] between these two parameters, which suggests that the observed temperature variations are correlated with occupancy patterns.
@@
Open data access allows project participants and external researchers to obtain data for further analysis from various perspectives. The data sharing application (available at https://historicbuildings.azurewebsites.net/download) enables users to choose various sampling intervals (raw or downsampled), parameters, and time periods. Downloaded data are cleaned of two types of anomalies. The first type is outliers, characterized by unexpected spikes or drops in the data, typically resulting from temporary sensor malfunctions or external interference. To address these outliers, threshold-based filtering was applied to remove data points that exceed predefined upper and lower bounds. The second type involves missing data, primarily due to communication issues. To fill in these missing data, linear interpolation was employed to estimate the missing values based on adjacent data points. Cleaned data are provided in comma-separated values (CSV) format, making it accessible for analysis with various tools.
Compared to existing digital twin platforms, these data applications offer greater customization and flexibility, allowing users to visualize specific environmental parameters and analyze data from selected sensor boxes within customizable date ranges. They provide real-time insights for quick identification of trends and correlations. Additionally, user-friendly interfaces simplify complex datasets, making them accessible to stakeholders with varying technical expertise, thus supporting effective management and decision-making for historic building conservation.
@@ IV-E Analysis of Indoor Climate
For question 2, indoor and outdoor humidity mixing ratios (MRs) over a year were compared to identify potential moisture sources. The MR is a measure of how much moisture is present in the air relative to the amount of dry air, expressed in grams of water vapor per kilogram of dry air (g/kg) [26]. This ratio provides a clear understanding of moisture content in the air. A comparison between indoor and outdoor MRs suggests a water source or sink in a building [27]. If the average indoor MR consistently exceeds the outdoor MR, it might indicate constant evaporation from the ground or walls of the building. MR is calculated by (1).
MR= $\text{MR}\ =\ 38.015\times\frac{10^{\frac{7.65t}{243.12+t}}\times\text{RH}}{p-\left(0.06112\times 10^{\frac{7.65t}{243.12+t}}\times\text{RH}\right)},$ (1)
where $t$ is temperature (C), RH is relative humidity (%), $p$ is atmospheric pressure (hPa). Since indoor atmospheric pressure was not measured, $p$ was substituted with a standard pressure of 1013 hPa for approximate calculations of indoor MRs, as recommended by the standard EN 16242:2012 [26]. Using the standard pressure as an approximation for indoor air pressure would slightly reduce the calculated indoor MRs. However, the resulting error is below 1%, which is within an acceptable range and does not impact comparisons between indoor and outdoor MRs. Outdoor MRs were calculated using collected atmospheric pressure.
@@
$\text{LIM}_{\text{I}}=\ \text{cosh}(0.128324\times(30-t))+75,$ (2)
where $t$ is temperature (C). The calculated value is expressed as a percentage (%). A RH value exceeding the LIMI at a given temperature is considered risky.
Short-term fluctuations of RH throughout a year were examined to determine the target range of RH and identify risks associated with strain-stress cycles, following the standard EN 15757:2010 [24]. Three key statistics representing the average levels and variability of RH were calculated: (1) the annual average RH, (2) seasonal cycles, and (3) short-term fluctuations. The annual average is the arithmetic mean of RH readings over one year. Seasonal cycles were determined by calculating a 30-day centered moving average (CMA), which computes the mean RH for each reading based on the 15 days preceding and following the measurement. Short-term fluctuations were calculated as the deviation between the current RH reading and the 30-day CMA, accounting for both natural seasonal variations and the stress-relaxation time constant of the materials. A safe band is determined by the 7th and 93rd percentiles of recorded RH fluctuations over the monitoring period, identifying the 14% of the most extreme and potentially risky fluctuations. If RH variations from the seasonal level are within 10%, a deviation of up to 10% from the seasonal RH level is considered acceptable as the safe band.
@@
The Pearson correlation coefficient ($r$) was used to measure the strength of a linear relationship between paired data. The magnitude of $r$ was classified as in study [64]:
  • $|r|<0.20$: very weak
  • $0.400.20\leq|r|<$ 0.40: weak
  • $0.600.40\leq|r|<$ 0.60: moderate
  • $0.80.60\leq|r|<$ 0.8: strong
  • $|r|\geq 0.80$: very strong
The Mann-Whitney U test [66] was used to assess differences between two groups of data, either in distribution or median. If the distributions differed in shape, the test evaluated differences in distribution. If the distributions had the same shape, it assessed differences in the median. A significance level of 0.05 was used. Statistical analysis was conducted using Python libraries pandas (v2.2.2) and SciPy (v1.13.1).
@@ V-A Temperature and Relative Humidity Throughout a Year
In rooms not equipped with heating systems, the trend of indoor temperature closely follows outdoor temperature. In the main building of Löfstad Castle, rooms in the basement and rooms on the ground and second floors lack heating systems. As shown in Fig. 9a and Fig. 9b, temperature in these rooms exhibited a similar trend to outdoor temperature throughout the year. For example, there was a very strong positive correlation between indoor temperature of Room 3 and the outdoor temperature ($r=0.84$). This suggests that the variations in indoor temperature in Room 3 were largely driven by fluctuations in outdoor temperature. Rooms on the first floor are equipped with heating systems. When these systems were turned off, the trend of indoor temperature was also driven by outdoor temperature. This pattern was evident from May to September 2023 (see Fig. 9b) for rooms on the first floor. However, when heating systems were working during the heating season, these rooms maintained a stable temperature around 16 C.
@@
Indoor temperature of different rooms on the same floor tend to be relatively close, while noticeable differences occur between rooms on different floors. During summer months, indoor temperature generally increases as a floor goes up. As depicted in Fig. 9b, around early July 2023, temperature in rooms on the first and second floors was approximately 25 C. Temperature of Room 222 on the second floor even reached up to 30 C. In contrast, during the same period, temperature of rooms on the ground floor and in the basement were approximately 20 C and 12 C, respectively (see Fig. 9a). The Mann-Whitney U test results indicated a significant difference in indoor temperature between rooms on different floors ($p<0.001$).
Abnormal variations in indoor temperature were observed during specific periods. As shown in Fig. 9a, Room 17 (without heating) on the ground floor experienced an unusually high temperature in the first half of December 2023. Moreover, heated rooms like Rooms 113 and 117 exhibited temperature spikes in November and December 2023 (see Fig. 9b). These anomalies may be attributed to maintenance activities conducted in these rooms.
@@
Rooms on the same floor typically have similar levels of RH. As depicted in Fig. 10a, RH in both basement rooms remained close to 100% throughout the year. This extremely high humidity is primarily due to the direct soil flooring, which allows groundwater to continuously evaporate from the wet soil into the indoor air. Likewise, Room 3 and Room 17 on the ground floor (see Fig. 10a), as well as the three rooms on both the first and second floors (see Fig. 10b), each have approximate levels of RH. An exception was observed in Room 5 on the ground floor (see Fig. 10a), where the sensor box detected higher RH compared to other rooms on the same floor. The RH frequently exceeded 80% and approached 100% from July to October 2023. The high RH readings are likely caused by the evaporation of water from the ground and moisture from the basement rise through walls.
There are notable differences in RH across floors. As illustrated in Fig. 10a and Fig. 10b, in rooms without heating systems (BF, GF, and 2F), RH decreased as the floor goes up. Conversely, in rooms with heating systems (1F), RH was lower compared to rooms on the second floor when the heating system was working. The Mann-Whitney U test results also indicated a significant difference in indoor RH between rooms on different floors ($p<0.001$).